A pitcher accelerates a 145g baseball from rest to 28.1 m/s. If this motion takes 0.27 seconds,

what average power was developed?
Answer #1

power = energy /time

energy = final KE - Initial kE

= 1/2 ( 145x 10^-3) ( 28.1^2) - 0= 57. 2467 J

power =  57. 2467 J / 0.27 = 212..025 watts apprx